OpenStudy (abhisar):

Does \(sin^2\omega t\) represents SHM?

imqwerty (imqwerty):

hmm haven't yet started SHM still on thermodynamics but yeah i know lil bit okay well like i have seem a lil ques which are like this- find the period of the function- \(y=sin (\omega t) \) .. like this are you talking about some term related to the topic which has that thing in it?

imqwerty (imqwerty):

its just a simple function which represents the Simple Harmonic motion

imqwerty (imqwerty):

yeah in that case \(sin^2 \omega t\) is a function representing the harmonic motion so we have - \(y=sin^2 \omega t\) to become simple harmonic \(\large \frac{d^2y}{dt^2} ∝ y\) here in our case \(\large\frac{d^2y}{dt^2}=2 \omega^2[cos^2 \omega t-sin^2 \omega t]=2\omega^2\left( 1-\frac{2y}{a}\right)\) yes it is directly proportional to y so its SHM

OpenStudy (abhisar):

O_O

imqwerty (imqwerty):

okay which part is troubling?

imqwerty (imqwerty):

oh wait i did a lil mistake

OpenStudy (abhisar):

Why have you taken -y and how do you know that for SHM double differentiation shoul dbe proportional to y

imqwerty (imqwerty):

in calculation part

OpenStudy (abhisar):

Yes.

imqwerty (imqwerty):

i added that a from no where..

imqwerty (imqwerty):

well.. i am not really sure about the reason :/ but well SHM's differential equation is this- \(\large\frac{d^y}{dt^2}=\omega^2 y\) so acc to this eq y must be directly proportion to the double derivative of the function for it to represent SHM

imqwerty (imqwerty):

*\(\frac{d^2 y}{dt^2}\)

OpenStudy (abhisar):

I am trying to understand how they have done in that book. Do you understand how they have done it?

OpenStudy (vincent-lyon.fr):

It is harmonic motion with non-zero rest position and \(2\omega\) angular frequency. \(sin^2(\omega t)=\frac12-\frac12\cos(2\omega t)=x_0+A\cos(\omega _0t+ \phi)\) with Rest position: \(x_0=\frac12\) Amplitude: \(A=\frac12\) Angular frequency: \(\omega_0 =2\omega\) Phase at t=0: \(\phi=\pi\)

OpenStudy (abhisar):

How \(sin^2(\omega t)=\frac12-\frac12\cos(2\omega t)\)

imqwerty (imqwerty):

we know that \(cos\theta =cos(\theta+2\pi)\) so if we have our equation like this-> \(x=Acos(\omega t +\alpha)\) then- \(x=Acos(\omega t + \alpha + 2 \pi) =A cos \left( \omega\left( t+\frac{2 \pi}{\omega} \right)+\alpha\right)\) lets say that \(t'=t+\frac{2\pi}{\omega}\) so \(x=Acos(\omega t'+\alpha)\) the this shows that the function at t coincides with the function at t'. and the interval (t'-t) is the period of x

imqwerty (imqwerty):

similarly in that link the equation was converted to this format so that we can find the time period easily

OpenStudy (abhisar):

Humse naa ho payega bhaiya ye ab xD

imqwerty (imqwerty):

lol xD

Parth (parthkohli):

What *is* SHM? That's the real question.

OpenStudy (abhisar):

Yes I think so..

OpenStudy (abhisar):

I don't know that quite clearly in first place.

Parth (parthkohli):

Q: What is SHM? A: Yes I think so... Full marks

OpenStudy (abhisar):

Any function that can be expressed in the form of Asin(wt + \(\sf \phi\)). Correct?

Parth (parthkohli):

That's not quite the definition of SHM, even though it is the solution to the equation.

OpenStudy (vincent-lyon.fr):

Definitely not.

OpenStudy (vincent-lyon.fr):

For instance: \(\sin^2(\omega t)\) can, but \(\sin^3(\omega t)\) can't.

Parth (parthkohli):

To check if anything is *SIMPLE* harmonic motion, check if its double-derivative is directly proportional to the position function. In other words, check if their ratio is constant. If it is, then yes.

OpenStudy (vincent-lyon.fr):

My question for the English-speaking friends here is: What does "simple" in simple harmonic motion refer to? In French, we just have "Oscillateur harmonique", with no hint that there should be a simple type and a non-simple one.

Parth (parthkohli):

\[x'' = -kx\]\[x = \sin^2 \omega t \]\[x' = 2\omega \sin \omega t \cos \omega t = \omega \sin (2\omega t) \]\[x'' = 2\omega^2 \cos (2\omega t)\]The required ratio is not constant. @Vincent-Lyon.Fr In India, simple harmonic motion is introduced as any oscillation that can be modelled by the above equation.

OpenStudy (vincent-lyon.fr):

Yes, but why not "Harmonic motion", full stop?

OpenStudy (vincent-lyon.fr):

In that case your initial function does not represent a SHM since the rest-position is not zero.

Parth (parthkohli):

Ah, I see what you're saying now. Nice.

Parth (parthkohli):

Good... I imagine you're right.

Parth (parthkohli):

However one wonders why the question is so ambiguous. There is no left-hand-side, i.e., saying something like \(x = \sin^2 \omega t\).

OpenStudy (vincent-lyon.fr):

I find this case strange, because although it can be meaningful in maths to distinguish between \(A\sin(\omega _0t+ \phi)\) and \(A\sin(\omega _0t+ \phi)+x_0\), in physics a simple change in the choice of origin does not change the nature of the motion itself.

Parth (parthkohli):

Well said. Didn't think of that one.

OpenStudy (vincent-lyon.fr):

What I call simple harmonic motion, is a function \(x(t)\) that can be solution of a differential equation: \(\ddot x(t)+\omega_0^2\;(x(t)-x_{rest})=0\) but the definition might be different in India, imposing \(x_{rest}=0\) ;-)

Parth (parthkohli):

It's obviously not different, lol. Just that in all contexts I've seen so far, the position function has been modeled in terms of the rest position, unless specified otherwise. I'll keep that one in mind though. Thank you very much.

OpenStudy (abhisar):

Can any one explain me How \(sin^2(\omega t)=\frac12-\frac12\cos(2\omega t)\)

Parth (parthkohli):

\[\cos (2\theta) = \cos^2 \theta - \sin^2 \theta = 1 - 2 \sin^2 \theta \]Now\[2 \sin^2 \theta = 1 - \cos (2 \theta)\]Divide both sides by two

Parth (parthkohli):

@Abhisar what's the answer given to you?

OpenStudy (vincent-lyon.fr):

Both following equations are very useful in physics: \(2 \cos^2 \theta = 1 +\cos (2 \theta)\) \(2 \sin^2 \theta = 1 - \cos (2 \theta)\)

OpenStudy (abhisar):

Thanks, I got it now. I don't have a background in maths so i was puzzled how you did that.

OpenStudy (abhisar):

Thank you very much for the time all of you guys :)

OpenStudy (irishboy123):

simple harmonic motion is motion pursuant to a **linear restorative** force. ie F = - kx. so \(m \ddot x + k x = 0\) with solution \(x(t) = A \cos (\omega t +\delta) \), same as \(x(t) = A \sin (\omega t +\delta) \) i am pretty sure simple means no driver and no damping.

OpenStudy (farcher):

y=sin^2(wt) is an oscillatory function but does not represent shm. If you know about differentiation then differentiate y with respect to time t twice and you will find that you do not get a result which is a constant times the original function. For example if y=sin(at) then the second differential is -w^2 sin(wt) = constant times y. Your function does not produce such a result so it is not a representation of shm.

OpenStudy (mathmate):

@vincent-lyon.fr You have a good point about HM and SHM. Guess "oscillateur harmonique" covers more cases than SHM. see also https://en.wikipedia.org/wiki/Complex_harmonic_motion

OpenStudy (abhisar):

@Vincent-Lyon.Fr , I think this is what you wanna trying to ask. Is it correct? 

OpenStudy (astrophysics):

https://www.youtube.com/watch?v=SZ541Luq4nE

OpenStudy (abhisar):

Awesome!

OpenStudy (vincent-lyon.fr):

@Abhisar IrishBoy123 gave the answer: 'simple' is opposed to 'driven' or 'damped' It need not be linear.